OpenStudy (trancenova):

Can anyone help me with stats?... Some researchers have developed a theory regarding the genetics determining resistance of a certain weed to a particular herbicide. According to this theory, one quarter of the plants raised in an experiment should be killed by the herbicide and three quarters should survive. When 50 plants are raised and sprayed with the herbicide, the researchers find that 32 of them survive and 18 die. Is the question, and I need to find the statistical test, test their hypothesis and..

OpenStudy (trancenova):

find the "95% confidence interval for the proportion of plants that will survive the herbicide in future experiments with large numbers of plants"

OpenStudy (trancenova):

I've used the Exact test for goodness-of-fit, is that correct? and I found that p = 0.0321, but not sure how to find the confidence interval or if I've used the right test.

OpenStudy (anonymous):

I was thinking about simple z-tests.

OpenStudy (trancenova):

Yeah that's what I thought at first but z-tests are based on averages, not count data :/

OpenStudy (anonymous):

Not sure what you mean by that. We have P(A) and P(A)', where A and A' are exhaustive and mutually exclusive. We have trial number n. We have μ=P(A).

OpenStudy (anonymous):

I may be speaking out of my behind here, but I'm tired, and I need to sleep. XD Good night, and good luck.

OpenStudy (trancenova):

lol thanks, goodnight

OpenStudy (dumbcow):

ahh i found it. for proportions the standard error is: \[\sqrt{\frac{p(1-p)}{n}}\] where p = 32/50 = 0.64, n = 50 margin of error(m) = z*SE confidence interval = p +- m for 95% choose z=1.96

OpenStudy (anonymous):

I'm back just to yell "binomial test" really loudly. Alright now, for reals, I'm going to sleep. XD

OpenStudy (trancenova):

hahahaha okay, binomial test does look right now :P